OpenStudy (anonymous):

solve for k 4(k+3)=(69-k)/2 HELP!!!

OpenStudy (anonymous):

4(k+3)=(69-k)/2 Multiply both sides by 2 to get 2 out of the denominator. 2(4)(k+3)=69-k 8(k+3) = 69-k Distribute 8 using distributive property 8k+24 = 69-k 8k+24+k = 69 9k = 69-24 k = 45/9 k=5

OpenStudy (anonymous):

Thanks so much do you mind helping with another question need to pass this class!!!??/

OpenStudy (anonymous):

just look them up